# Unified Modeling Language Code Generation from Diagram Images Using Multimodal Large Language Models

### v1

@@ ABSTRACT
The Unified Modeling Language is a standardized visual language widely used for modeling and documenting the design of software systems. Although many tools generate UML diagrams from UML code, generating executable UML code from image-based UML diagrams remains challenging. This paper proposes a new approach to generate UML code using a large multimodal language model automatically. Synthetic UML activity and sequence diagram datasets were created to train and test the model. We compared standard fine-tuning with LoRA techniques to optimize base models. The experiments measured code generation accuracy across different model sizes and training strategies. These results demonstrated that domain-adapted MM-LLMs perform for UML code generation automation, whereby, at the best model, it achieved BLEU and SSIM scores of 0.779 and 0.942 on sequence diagrams. This will enable the modernization of legacy systems and decrease the manual effort in software development workflows.
Keywords: UML, Large Language Models, Machine Learning, Code Generation
@@ 1 Introduction
In software engineering, the Unified Modeling Language (UML) is widely used to represent and analyze complex systems visually [2]. UML provides a standardized framework for documenting, designing, and refining software solutions [19, 20]. It supports various stages of the development lifecycle by structuring software design and clarifying relationships among components, encouraging collaboration among stakeholders [11]. Metzner compares UML proficiency to an architect’s use of blueprints, highlighting its importance in creating systematic, organized design frameworks [20].
@@ 1.1 Addressing the Problem
Automated code generation from visual representations has gained traction in recent years. For example, pix2code demonstrated the conversion of GUI screenshots into executable code across platforms like iOS, Android, and the web [1]. Using CNNs and LSTM networks, it bypasses complex heuristics by interpreting GUI elements directly from images. Building on this work, Cai et al. [3]’s GUICG enhances component localization and classification through CNN-based target detection, achieving notable accuracy improvements on large-scale GUI datasets. Beyond GUIs, Ellis et al. explored code generation from freehand sketches using a two-stage model combining CNNs with program synthesis techniques [9]. Identifying geometric primitives and constructing structured code representations highlights the potential of machine learning in bridging visual and code-based workflows. Researchers have also started developing models that extract elements from UML diagrams stored in non-editable formats. For instance, ReSECDI identifies core UML components like classes and relationships across varying formats [4]. However, while ReSECDI provides syntactic extraction, it does not yet generate executable code, underscoring the need for machine learning techniques to bridge this gap [7].
Large Language Models (LLMs) have demonstrated impressive capabilities in understanding and generating human-like text. Advanced transformer-based architectures such as BERT [8], GPT [24], and T5 [25] capture contextual relationships within text, enabling tasks like summarization, question answering, and code generation. However, LLMs process only textual data, limiting their utility in domains requiring multimodal understanding, such as image or video analysis. Researchers developed MM-LLMs that incorporate visual, auditory, and textual data into unified frameworks to address these limitations. Models like CLIP [23] align visual embeddings with textual embeddings, enabling cross-modal tasks such as visual question answering (VQA) and zero-shot image classification. For example, CLIP (Contrastive Language–Image Pretraining) learns shared visual and textual semantics from large-scale datasets of image-text pairs, effectively bridging linguistic context and visual perception. MM-LLMs leverage these multimodal capabilities to tackle tasks that traditional LLMs cannot address, such as interpreting UML diagrams and generating corresponding executable code. MM-LLMs align diagram structures with text-based outputs by combining visual encoders with language models, enhancing accuracy and flexibility for applications like UML-to-code automation.
@@ 1.4 Automating Code Generation
LLMs have significantly advanced automated code generation, providing tools that analyze textual and structural information to generate functional code. For example, Meta’s CodeCompose offers developers real-time code suggestions and streamlining routine coding tasks through context-aware predictions [21]. Similarly, OpenAI’s Codex, fine-tuned from GPT-3 on GitHub datasets, achieves a notable success rate in solving coding challenges through iterative sampling strategies [5]. These tools demonstrate the potential of LLMs to improve productivity by reducing manual intervention in code generation. Despite their success, traditional LLMs are limited in diagram-to-code automation. Converting UML diagrams into executable code demands both linguistic and visual comprehension. Thus, the use of MM-LLMs fine-tuning techniques play a critical role in enhancing MM-LLMs for UML-to-code generation. Instruction-based fine-tuning and Chain-of-Thought prompting improve the models’ ability to generate contextually accurate and logically consistent code. Recent studies have shown that fine-tuned models generalize effectively across benchmarks, maintaining structural fidelity in generated outputs [6]. Studies like Plot2Code and Design2Code highlight the challenges of converting structured visuals into functional code. For instance, while GPT-4 achieved high accuracy in generating executable code from scientific plots, its performance declined with increasing visual complexity [29, 32]. These findings underscore the importance of tuning MM-LLMs to handle intricate UML diagrams, where maintaining relational and sequential accuracy is critical. In this research, we extend these advancements by fine-tuning MM-LLMs to generate code from activity and sequence diagrams autonomously. The models are trained to align diagram structures with functional outputs, ensuring high syntactic and structural fidelity.
@@ 2 Methodology
Automating UML code generation requires models capable of processing visual and textual data. We use the Large Language and Vision Assistant (LLaVA) [15], specifically its improved version, LLaVA-1.5 [13], a foundational model. Both LLaVA and LLaVA-1.5 integrate a CLIP-based visual encoder [18] with the Vicuna language model [17] via a vision-language connector, enabling the extraction of visual features from UML diagrams and translating them into accurate textual outputs. LLaVA uses a linear projection layer to align visual embeddings with the language model. In contrast, LLaVA-1.5 introduces a multi-layer perceptron as a more complex connector, enabling nuanced interactions between visual and textual representations. The architectural enhancements, combined with expanded visual instruction tuning on VQA and diagram datasets, improve LLaVA-1.5’s performance for tasks requiring detailed diagram analysis, such as UML-to-code generation.
@@ 2.1 Fine-tuning and LoRA Fine-tuning
LoRA reduces the number of trainable parameters by introducing trainable low-rank matrices into the architecture. Specifically, instead of updating all parameters in the model, LoRA modifies only a subset of the parameters in the form of low-rank matrices, significantly decreasing memory consumption during training. The low-rank approximation minimizes the number of trainable parameters by focusing only on updating specific components, leading to substantial reductions in memory usage and computational overhead [10]. Despite this simplification, LoRA can demonstrate performance comparable to full fine-tuning in several large-scale language models. Given the scale of models like LLaVA-1.5, incorporating LoRA enables the system to maintain computational efficiency without compromising its ability to adapt to the unique requirements of tasks such as UML diagram analysis. This balance between resource efficiency and adaptability makes LoRA a compelling approach for fine-tuning large models on domain-specific tasks.
@@ 2.2 Hyperparameters
The hyperparameters lora_r and lora_alpha are central to LoRA fine-tuning. lora_r determines the rank of the low-rank matrices integrated into the model. Increasing lora_r enhances the expressive power of these matrices, enabling more nuanced adaptations but requiring greater memory and computational resources [10]. lora_alpha, a scaling factor, adjusts the influence of the low-rank matrices during training. Increasing lora_alpha amplifies the impact of the matrices on the model output, enhancing their prominence in the learning process. The mm_projector_lr hyperparameter sets the learning rate for the projector layer. Since LoRA fine-tuning updates only a subset of parameter values, the projector learning rate ensures balanced updates concerning the rest of the model. LoRA fine-tuning employs a higher learning rate and larger batch size, as updating fewer parameters than standard fine-tuning allows the model to converge more efficiently. A batch size of 8 was used for standard fine-tuning, while LoRA fine-tuning utilized a batch size of 16 to accommodate its more efficient parameter updates. Despite the differences in batch size, the training times and computational performance metrics provide valuable insights into each approach’s efficiency and scalability.
@@ 2.3 Synthetic Data Generation
Two types of synthetic UML diagrams, activity and sequence diagrams, were generated using PlantUML, a text-based tool to create visual diagrams. Randomized text strings based on PlantUML syntax created diverse and scalable datasets that effectively mimic UML diagrams. PlantUML syntax facilitates the direct translation of UML concepts into textual form, enabling automated rendering into PNG or SVG images. The sequence diagrams emphasize participant interactions and message flows, as illustrated in Fig 5. The corresponding PlantUML code showcases the structured textual representation of visual relationships.
Similarly, the activity diagrams depict processes with decisions and concurrent actions, represented visually in Fig. S1 and in PlantUML text format (Fig. S2). These paired examples demonstrate the translation between visual UML representations and their underlying textual descriptions, essential for automated code generation. The diagrams and corresponding code files were systematically saved with unique identifiers to ensure dataset traceability and prevent duplication. Fig. S3 provides an example of the JSON data format used during model training, aligning each diagram image with its UML code.
@@ 2.3.1 Dataset Breakdown
The dataset is then split into testing and training with an 80-20 split. The generation pipeline uses a Python script to tokenize the PlantUML code, align it with input images, and store the data in JSON format for training. The testing split had an additional step where data was organized in a question-answer format and added as another JSON file. The structured format supports effective multimodal learning, as demonstrated below.
@@ 2.4 Evaluation Metrics
The BLEU metric, introduced by Papineni et al. [22], is widely used for evaluating text generation tasks. It measures the similarity between generated text and a reference by comparing n-grams, offering a quantitative measure of syntactic accuracy. BLEU’s simplicity, scalability, and independence from specific programming languages make it well-suited for large-scale evaluation of generated UML code. In prior studies, BLEU has been successfully applied to similar tasks such as evaluating large-scale language models for code generation [5, 33]. A higher BLEU score indicates stronger alignment between the generated output and the ground truth, ensuring that syntactic fidelity is maintained during UML-to-code translation.
The Structural Similarity Index Measure (SSIM) is a common image quality metric that evaluates the visual similarity between generated and reference images [31, 27]. Unlike pixel-based metrics, SSIM considers luminance, contrast, and structural features to align with human visual perception. This makes SSIM especially relevant for UML diagrams, as it prioritizes the layout and relationships between elements over minor pixel variations. SSIM produces scores ranging from -1 to 1, where higher values reflect closer visual resemblance. When evaluating UML diagrams, SSIM helps ensure that structural elements, such as object relationships, decision nodes, and component alignment, remain intact. Figs S4 and S5 demonstrate a high-scoring diagram, ground truth and generated, respectively. Figs S6 and S7 demonstrate a low-scoring diagram, ground truth and generated, respectively. A high SSIM and BLEU score signifies that the generated UML diagram aligns visually and textually with the reference, preserving spatial arrangement and semantic content. Conversely, low scores indicate discrepancies, such as misaligned components or incorrect labels, compromising structural fidelity. These examples underscore the critical role of SSIM and BLEU in validating the performance of models tasked with UML code generation. Maintaining high visual and textual accuracy is essential to ensure the diagrams remain functional for software development workflows.
@@
The error analysis follows a structured evaluation process inspired by methodologies from prior studies [7]. Three main error types are examined: syntax errors, UML code absence, and diagram mismatches. Syntax errors refer to missing symbols or incomplete code blocks that prevent the diagram from rendering correctly. These errors are flagged automatically when the PlantUML rendering tool fails to generate valid diagrams. UML absence errors occur when the generated output is irrelevant or incomplete and contains placeholder text or non-UML responses. The diagram mismatch score quantifies logical inconsistencies, such as producing an incorrect diagram type (e.g., a class diagram instead of a sequence diagram). These error metrics comprehensively evaluate the model’s reliability, with lower scores indicating better performance.

### v2

@@ Abstract
The Unified Modeling Language is a standardized visual language widely used for modeling and documenting the design of software systems. Although many tools are available that generate UML diagrams from UML code, generating executable UML code from image-based UML diagrams remains challenging. This paper proposes a new approach to generate UML code using a large multimodal language model automatically. Synthetic UML activity and sequence diagram datasets were created to train and test the model. We compared the standard fine-tuning with LoRA techniques to optimize base models. The experiments measured the code generation accuracy across different model sizes and training strategies. These results demonstrated that domain-adapted MM-LLMs perform for UML code generation automation, whereby, at the best model, it achieved BLEU and SSIM of 0.779 and 0.942 on sequence diagrams. This will enable the modernization of legacy systems and decrease the manual effort put into software development workflows.
keywords:
UML , Large Language Models , Machine Learning , Code Generation
keywords:
UML , Large Language Models , Machine Learning, Code Generation
journal: Machine Learning with Applicationsjournal: Machine Learning with Applications
[…]
[ou]organization=School of Computer Science, University of Oklahoma, addressline=110 W. Boyd St., city=Norman, state=OK, country=US
[…]
[ml]organization=Enterprise Architecture, MapLarge, addressline=1201 Peachtree Street NE, Building 400, Suite 1750, city=Atlanta, state=GA, country=US
[…]
[fsu]organization=School of Communication, Florida State University, addressline=4100 University Center, Building C, city=Tallahassee, state=FL, country=US
footnotetext: This is the accepted manuscript version of the article published in Machine Learning with Applications, Volume 20, 2025, Article 100660. The final version is available at https://doi.org/10.1016/j.mlwa.2025.100660.
@@ 1 Introduction
In software engineering, the Unified Modeling Language (UML) is widely used to represent and analyze complex systems visually (Booch et al., 1998). UML provides a standardized framework for documenting, designing, and refining software solutions (Lu and Alexandru, 2023; Metzner, 2024). It supports various stages of the development lifecycle by structuring software design and clarifying relationships among components, encouraging collaboration among stakeholders (Koç et al., 2021). Metzner compares UML proficiency to an architect’s use of blueprints, highlighting its importance in creating systematic, organized design frameworks (Metzner, 2024).
@@ 1.1 Addressing the Problem
Automated code generation from visual representations has gained traction in recent years. For example, pix2code demonstrated the conversion of GUI screenshots into executable code across platforms like iOS, Android, and the web (Beltramelli, 2018). Using CNNs and LSTM networks, it bypasses complex heuristics by interpreting GUI elements directly from images. Building on this work, Cai et al. (2023)’s GUICG enhances component localization and classification through CNN-based target detection, achieving notable accuracy improvements on large-scale GUI datasets. Beyond GUIs, Ellis et al. explored code generation from freehand sketches using a two-stage model combining CNNs with program synthesis techniques (Ellis et al., 2018). Identifying geometric primitives and constructing structured code representations highlights the potential of machine learning in bridging visual and code-based workflows. Researchers have also started developing models that extract elements from UML diagrams stored in non-editable formats. For instance, ReSECDI identifies core UML components like classes and relationships across varying formats (Chen et al., 2022). However, while ReSECDI provides syntactic extraction, it does not yet generate executable code, underscoring the need for machine learning techniques to bridge this gap (Conrardy and Cabot, 2024).
Large Language Models (LLMs) have demonstrated impressive capabilities in understanding and generating human-like text. Advanced transformer-based architectures such as BERT (Devlin et al., 2018), GPT (Radford et al., 2018), and T5 (Raffel et al., 2020) capture contextual relationships within text, enabling tasks like summarization, question answering, and code generation. However, LLMs process only textual data, limiting their utility in domains requiring multimodal understanding, such as image or video analysis. Researchers developed MM-LLMs that incorporate visual, auditory, and textual data into unified frameworks to address these limitations. Models like CLIP (Radford et al., 2021) align visual embeddings with textual embeddings, enabling cross-modal tasks such as visual question answering (VQA) and zero-shot image classification. For example, CLIP (Contrastive Language–Image Pretraining) learns shared visual and textual semantics from large-scale datasets of image-text pairs, effectively bridging linguistic context and visual perception. MM-LLMs leverage these multimodal capabilities to tackle tasks that traditional LLMs cannot address, such as interpreting UML diagrams and generating corresponding executable code. MM-LLMs align diagram structures with text-based outputs by combining visual encoders with language models, enhancing accuracy and flexibility for applications like UML-to-code automation.
@@ 1.4 Automating Code Generation
LLMs have significantly advanced automated code generation, providing tools that analyze textual and structural information to generate functional code. For example, Meta’s CodeCompose offers developers real-time code suggestions and streamlining routine coding tasks through context-aware predictions (Murali et al., 2024). Similarly, OpenAI’s Codex, fine-tuned from GPT-3 on GitHub datasets, achieves a notable success rate in solving coding challenges through iterative sampling strategies (Chen et al., 2021). These tools demonstrate the potential of LLMs to improve productivity by reducing manual intervention in code generation. Despite their success, traditional LLMs are limited in diagram-to-code automation. Converting UML diagrams into executable code demands both linguistic and visual comprehension. Thus, the use of MM-LLMs fine-tuning techniques play a critical role in enhancing MM-LLMs for UML-to-code generation. Instruction-based fine-tuning and Chain-of-Thought prompting improve the models’ ability to generate contextually accurate and logically consistent code. Recent studies have shown that fine-tuned models generalize effectively across benchmarks, maintaining structural fidelity in generated outputs (Chung et al., 2022). Studies like Plot2Code and Design2Code highlight the challenges of converting structured visuals into functional code. For instance, while GPT-4 achieved high accuracy in generating executable code from scientific plots, its performance declined with increasing visual complexity (Shi et al., 2024; Wu et al., 2024). These findings underscore the importance of tuning MM-LLMs to handle intricate UML diagrams, where maintaining relational and sequential accuracy is critical. In this research, we extend these advancements by fine-tuning MM-LLMs to generate code from activity and sequence diagrams autonomously. The models are trained to align diagram structures with functional outputs, ensuring high syntactic and structural fidelity.
@@ 2 Methodology
Automating UML code generation requires models capable of processing visual and textual data. We use the Large Language and Vision Assistant (LLaVA) (Liu et al., 2024c), specifically its improved version, LLaVA-1.5 (Liu et al., 2024a), a foundational model. Both LLaVA and LLaVA-1.5 integrate a CLIP-based visual encoder (Lu et al., 2024) with the Vicuna language model (LMSYS, 2023) via a vision-language connector, enabling the extraction of visual features from UML diagrams and translating them into accurate textual outputs. LLaVA uses a linear projection layer to align visual embeddings with the language model. In contrast, LLaVA-1.5 introduces a multi-layer perceptron as a more complex connector, enabling nuanced interactions between visual and textual representations. The architectural enhancements, combined with expanded visual instruction tuning on VQA and diagram datasets, improve LLaVA-1.5’s performance for tasks requiring detailed diagram analysis, such as UML-to-code generation.
@@ 2.1 Fine-tuning and LoRA Fine-tuning
LoRA reduces the number of trainable parameters by introducing trainable low-rank matrices into the architecture. Specifically, instead of updating all parameters in the model, LoRA modifies only a subset of the parameters in the form of low-rank matrices, significantly decreasing memory consumption during training. The low-rank approximation minimizes the number of trainable parameters by focusing only on updating specific components, leading to substantial reductions in memory usage and computational overhead (Hu et al., 2021). Despite this simplification, LoRA can demonstrate performance comparable to full fine-tuning in several large-scale language models. Given the scale of models like LLaVA-1.5, incorporating LoRA enables the system to maintain computational efficiency without compromising its ability to adapt to the unique requirements of tasks such as UML diagram analysis. This balance between resource efficiency and adaptability makes LoRA a compelling approach for fine-tuning large models on domain-specific tasks.
@@ 2.2 Hyperparameters
The hyperparameters lora_r and lora_alpha are central to LoRA fine-tuning. lora_r determines the rank of the low-rank matrices integrated into the model. Increasing lora_r enhances the expressive power of these matrices, enabling more nuanced adaptations but requiring greater memory and computational resources (Hu et al., 2021). lora_alpha, a scaling factor, adjusts the influence of the low-rank matrices during training. Increasing lora_alpha amplifies the impact of the matrices on the model output, enhancing their prominence in the learning process. The mm_projector_lr hyperparameter sets the learning rate for the projector layer. Since LoRA fine-tuning updates only a subset of parameter values, the projector learning rate ensures balanced updates concerning the rest of the model. LoRA fine-tuning employs a higher learning rate and larger batch size, as updating fewer parameters than standard fine-tuning allows the model to converge more efficiently. A batch size of 8 was used for standard fine-tuning, while LoRA fine-tuning utilized a batch size of 16 to accommodate its more efficient parameter updates. Despite the differences in batch size, the training times and computational performance metrics provide valuable insights into each approach’s efficiency and scalability.
@@ 2.3 Synthetic Data Generation
Two types of synthetic UML diagrams—activity and sequence diagrams—were programmatically generated using PlantUML, a text-based diagramming tool widely adopted for its ease of integration and automation. The diagrams were constructed using randomized text strings drawn from a curated vocabulary of the 25,286 most popular words found in English. These strings were inserted into standard PlantUML syntax structures, thereby generating a wide variety of visually valid UML diagrams that are structurally diverse but semantically nonsensical. Importantly, the diagrams do not follow any coherent application logic; rather, they serve as syntactic approximations of real UML diagrams. This approach was chosen to isolate the visual parsing and language modeling capabilities of the system without introducing domain-specific priors.
PlantUML enables direct conversion of textual UML descriptions into graphical formats, supporting both rasterized PNG and vectorized SVG outputs. Each synthetic diagram was exported in PNG format. The sequence diagrams emphasize participant interactions and message flows, as illustrated in Fig 5. The corresponding PlantUML code showcases the structured textual representation of visual relationships in Fig 6. In a similar fashion, activity diagrams were synthesized to reflect control flows including branching, merging, and parallel operations, represented visually in Fig. S1 and in PlantUML text format (Fig. S2). Their labels and process steps are composed of randomly selected words, resulting in workflows that lack meaningful interpretation. Nevertheless, they adhere to the syntactic rules of UML activity diagramming, making them suitable for testing model robustness to surface-level structural variation.
@@ 2.3.1 Dataset Breakdown
All generated diagrams and corresponding PlantUML source files were stored with unique identifiers to ensure dataset traceability and to prevent unintentional duplication. Additionally, each image-code pair was serialized into a JSON format (see Supplemental Fig. S3), which was used as the primary input format during model training. The dataset was partitioned using a simple random sampling procedure to produce an 80/20 train-test split. Testing examples were selected without regard to structure or content, ensuring unbiased evaluation across the synthetically generated samples. The testing split had an additional step where data was organized in a question-answer format and added as another JSON file. The structured format supports effective multimodal learning, as demonstrated below.
2.3.2 Real-World Testing Evaluation
While the majority of our dataset is synthetically generated to ensure controlled variability and sufficient volume for training, it is important to assess how well models perform on naturally occurring data. To this end, we curated a small set of real-world UML diagrams—29 activity diagrams and 28 sequence diagrams—totaling 57 examples. These were sourced from publicly available resources and serve as a benchmark for evaluating the generalization capabilities of models trained exclusively on synthetic data. The real-world data is used solely for testing and enables comparison across models with varying levels of training exposure. We evaluate the original model, a minimally trained small LoRA variant, two standard models (7B and 13B) trained on the small dataset, and finally, models trained extensively on the extra-large dataset using both LoRA and full parameter tuning. This setup allows us to examine the influence of training intensity—from no training to full-scale training—on the model’s ability to generalize to naturally structured UML diagrams. Including this evaluation provides valuable insights into the robustness and practical applicability of our models. It helps determine whether large-scale synthetic training introduces limitations or enhances the model’s ability to interpret real-world inputs. The distribution of the real-world testing dataset is summarized in Table 3.
The real-world UML diagrams were sourced from publicly available, educational, and technical resources that include software engineering tutorials, documentation repositories, and open-source guides. Specifically, we collected data from three primary sources: the official PlantUML guide (PlantUML, 2025), as well as tutorial-based resources on activity and sequence diagrams from WebDevTutor (Web Dev Tutor, 2023b, a). Selection criteria focused on clarity, diversity of structure, and representational fidelity to practical use cases in software design. All chosen diagrams were fully rendered and accompanied by the corresponding PlantUML source code, ensuring their suitability for both image-based and code-based model evaluation. Unlike the synthetic diagrams, which were generated by inserting randomized text strings into grammatically correct UML syntax templates, the real-world diagrams were authored by domain experts or educators to illustrate concrete workflows and interactions. These real diagrams typically exhibit meaningful semantic content, coherent control or message flows, and intentional labeling conventions that reflect real application logic or business processes. For instance, activity diagrams in the real-world set often include process names suggesting a clear operational sequence. Sequence diagrams show realistic inter-object communications such as login procedures or database queries, in contrast to the semantically nonsensical but syntactically valid interactions found in synthetic examples.
Table 3: Real-World Testing Data Distribution
Data Type Testing Size
Activity Diagrams 29
Sequence Diagrams 28
2.4 Hardware and Pipeline Devoplment
We completed all fine-tuning and evaluation tasks on the University of Oklahoma’s supercomputing infrastructure, leveraging an SLURM-managed HPC cluster. Depending on node availability, training was conducted on NVIDIA A100 GPUs, either 40GB or 80GB versions. For standard full-parameter fine-tuning—required for larger models like Vicuna-13B and LLaVA-13B—we utilized four A100 GPUs to meet the memory and compute demands. In comparison, LoRA fine-tuning was successfully executed using just two A100 GPUs. All jobs were submitted via SLURM scripts, which specified GPU resources, runtime configurations, and distributed training parameters.
Our approach followed the official LLaVA training pipeline from the project’s GitHub repository, incorporating full model and LoRA-based fine-tuning scripts. The process involved two key stages: feature alignment and visual instruction tuning. Depending on the experiment, we used either the standard (fine-tune.sh) or LoRA (finetune_lora.sh) script, with only minimal modifications tailored to our HPC setup and resource allocations (Liu et al., 2024a) Code available at: https://github.com/haotian-liu/LLaVA. Model checkpoints were saved at predefined intervals, and SLURM output was used to monitor logs in real-time. We performed all inference runs on a single A100 GPU per model for consistent evaluation, applying this to both our trained checkpoints and baseline models. The setup ensured uniform memory and compute conditions across all experiments. All associated pre- and post-processing—such as JSON parsing and results aggregation—was handled through Python 3.11 scripts within a Linux terminal environment.
@@ 2.5 Evaluation Metrics
The BLEU metric, introduced by Papineni et al. (Papineni et al., 2002), is widely used for evaluating text generation tasks. It measures the similarity between generated text and a reference by comparing n-grams, offering a quantitative measure of syntactic accuracy. BLEU’s simplicity, scalability, and independence from specific programming languages make it well-suited for large-scale evaluation of generated UML code. In prior studies, BLEU has been successfully applied to similar tasks such as evaluating large-scale language models for code generation (Chen et al., 2021; Zhang et al., 2023). A higher BLEU score indicates stronger alignment between the generated output and the ground truth, ensuring that syntactic fidelity is maintained during UML-to-code translation.
The Structural Similarity Index Measure (SSIM) is a common image quality metric that evaluates the visual similarity between generated and reference images (Wang et al., 2004; Reznik, 2023). Unlike pixel-based metrics, SSIM considers luminance, contrast, and structural features to align with human visual perception. This makes SSIM especially relevant for UML diagrams, as it prioritizes the layout and relationships between elements over minor pixel variations. SSIM produces scores ranging from -1 to 1, where higher values reflect closer visual resemblance. When evaluating UML diagrams, SSIM helps ensure that structural elements, such as object relationships, decision nodes, and component alignment, remain intact. Figs S4 and S5 demonstrate a high-scoring diagram, ground truth and generated, respectively. Figs S6 and S7 demonstrate a low-scoring diagram, ground truth and generated, respectively. A high SSIM and BLEU score signifies that the generated UML diagram aligns visually and textually with the reference, preserving spatial arrangement and semantic content. Conversely, low scores indicate discrepancies, such as misaligned components or incorrect labels, compromising structural fidelity. These examples underscore the critical role of SSIM and BLEU in validating the performance of models tasked with UML code generation. Maintaining high visual and textual accuracy is essential to ensure the diagrams remain functional for software development workflows.
@@
The error analysis follows a structured evaluation process inspired by methodologies from prior studies (Conrardy and Cabot, 2024). Three main error types are examined: syntax errors, UML code absence, and diagram mismatches. Syntax errors refer to missing symbols or incomplete code blocks that prevent the diagram from rendering correctly. These errors are flagged automatically when the PlantUML rendering tool fails to generate valid diagrams. UML absence errors occur when the generated output is irrelevant or incomplete and contains placeholder text or non-UML responses. The diagram mismatch score quantifies logical inconsistencies, such as producing an incorrect diagram type (e.g., a class diagram instead of a sequence diagram). These error metrics comprehensively evaluate the model’s reliability, with lower scores indicating better performance.